1091: Maintenance Goudsmit
One of the systemdisks reported SMART-errors and will be replaced.
CPK messages are initially sent to the CPK mailing list, you can (un)subscribe via this link. You can also follow the service interruption messages via RSS using the link in the title under the RSS icon. If the CPK takes more time to resolve, any updates are published on this website.
For RU wide service interruption see meldingen.ru.nl.
One of the systemdisks reported SMART-errors and will be replaced.
Lilo did not shut down correctly during the planned weekly reboot. After a hardware reset, the machine booted just fine
Just like yesterday, a reboot of the machine was necessary to restore the functionality. Tomorrow morning, the server will restart with a new kernel. It this doesn’t solve the problem, we will look into a new version of Samba.
Just like yesterday, a reboot of the machine was necessary to restore the functionality. We will investigate if we can remediate this problem with newer software versions (Samba, Ubuntu, …)
Some Eduroam users didn’t get an IP number, because one of the IP ranges had no free leases. This ranges contained all @science.ru.nl users. The ISC networking department first decreased the lease time. Now they also make this range free for use by @science.ru.nl users. This will prevent the problem from occuring again soon.
Again a reboot of the machine was necessary to restore the functionality. As a patch we will reboot this server each night.
The new RU IdM system erroneously rewrote the RU password hash when data about a person changed. This primariliy affected employees of FNWI, because for this group the Science email address had been incorporated in IdM. This will make it possible to use the Science address as an external mail address when resetting RU passwords. When the problem was recognized, it was fixed swiftly by restoring the RU password hashes from backup. ...
April 28/29/30, the Peage MFP and the Peage POS near the restaurant in the Huygens building cannot be used. The ISC will then investigate whether Peage can also be made available for employees instead of only for students. April 28, the MFP will be moved to the test location, April 30 it will be moved back.
The printer has to be repaired. Spare parts have been ordered. We hope to have the printer operational again mid-May.
Although the hardware of the server was replaced, the server didn’t react for unknown reasons. A reboot of the machine solved the problem. We will think about ways to tackle this problem.